X-ray from a certain cobalt target tube are composed of the strong ๐พ series


of Cobalt and weak ๐พ lines due to impurities. The wavelengths of ๐พ๐›ผ are


1.785๐ด


๐‘œ


for Cobalt and 2.285๐ด


๐‘œ


and 1.537๐ด


๐‘œ


for impurities. (a) What


elements are they? [For the ๐พ-series, a=1 in Moseleyโ€™s law]

1
Expert's answer
2021-12-10T12:30:19-0500

Moseley law

"\sqrt{\nu}=z-1"

"{\nu}=\frac{c}{\lambda}"

"\sqrt\frac{c}{\lambda}=z-1"

For cobalt "\lambda=\lambda _{co}"

impurities

"\lambda=\lambda_x"

"\sqrt{\frac{\lambda_{co}}{\lambda_x}}=\frac{z_x-1}{z_{co}-1}"

First impurities

"\lambda_{co}=1.785A°\\\lambda x_1=2.285A°\\\lambda x_2=1.537A°"

"\sqrt{\frac{\lambda_{co}}{{\lambda_x}_1}}=\frac{z_{x1}-1}{z_{co}-1}"

"\sqrt{\frac{1.785}{2.285}}=\frac{z_{x1}-1}{27-1}"

Both take square

"\sqrt{\frac{1.785}{2.285}}=\frac{z_{x1}-1}{27-1}"

"0.8838=\frac{z_{x1}-1}{26}"

"z_{x1}=23+1=24"

Z=24

Cromium



"\sqrt{\frac{\lambda_{co}}{\lambda_{x2}}}=\frac{z_{x2}-1}{z_{co}-1}"

Put value

"\sqrt{\frac{1.785}{1.537}}=\frac{z_{x2}-1}{27-1}"

"1.07766=\frac{z_{x2}-1}{26}"

"z_{x2}=28+1=29"

Z=29

copper